# Interpretable Plant Leaf Disease Detection Using Attention-Enhanced CNN

### v1

@@ 3.2.2 Spatial Attention Module
Fig. 5 illustrates the spatial attention module integrated into the proposed CBAM-VGG16 architecture. This module generates a 2D spatial attention map $S_{m}\in\mathbb{R}^{c\times W\times H}$, which is applied to the channel-refined feature map $I_{c}f^{\prime}$. To compute spatial attention, the module first applies average pooling and max pooling operations along the channel axis, resulting in two 2D feature maps: $I_{s}^{\text{avg}}\in\mathbb{R}^{1\times W\times H}$ and $I_{s}^{\text{max}}\in\mathbb{R}^{1\times W\times H}$. These two maps are then concatenated along the channel dimension and passed through a convolutional layer to produce the spatial attention map. The resulting map is used to refine the features by performing an element-wise multiplication with $I_{c}f^{\prime}$, yielding the final spatially-refined output $I_{s}f^{\prime\prime}$. The spatial attention computation in CBAM is formally described in Eq. 5 and  6.
$S_{m}(I_{c}f^{\prime})$ $=\sigma\big(\nu_{7\times 7}([\text{AvgPool}(I_{c}f^{\prime});\text{MaxPool}(I_{c}f^{\prime})])\big)$ (5)
$S_{m}(I_{c}f^{\prime})$ $=\sigma\big(\nu_{7\times 7}([I_{s}^{\text{avg}};I_{s}^{\text{max}}])\big)$ (6)
In the above equations, $\sigma$ represents the sigmoid activation function, and $\nu_{7\times 7}$ denotes a convolution operation with a kernel size of $7\times 7$. The spatial attention module computes a 2D attention map by first applying average pooling and max pooling operations along the channel dimension of the input $I_{c}f^{\prime}$, resulting in two spatial descriptors $I_{s}^{\text{avg}}$ and $I_{s}^{\text{max}}$. These are concatenated and passed through the convolution layer to obtain the spatial attention map $S_{m}(I_{c}f^{\prime})$, which is then element-wise multiplied with the input $I_{c}f^{\prime}$ to produce the final spatially-refined feature map $I_{s}f^{\prime\prime}$.
The intermediate layers of the model utilize the Rectified Linear Unit (ReLU) activation function [nair2010rectified], defined as:
@@
Figure 4: Channel attention module used in CBAM layers.
@@
The final output layer employs the Softmax activation function [Xiao], a normalized exponential function commonly used for multi-class classification. Softmax transforms the input vector into a probability distribution where the sum of all output probabilities is one, as defined in Eq. 8.
Figure 5: Spatial attention module used in CBAM layers
@@
In our work, as indicated by Eq. 9, we have utilised the Cross Entropy loss function [mao2023crossentropyl], for model optimization.
@@
Our proposed architecture have five CBAM layers which makes it a complex model and can cause overall classification of plant leaf disease detection to become less generic. To reduce the model’s complexity, L2 regularization [YANG2023118873] is employed where extreme changes in weights during the training phase are penalized by regularization reducing the probability of over-fitting. During training, the L2 regularizer additionally simplifies the input features and stabilises performance. The definition of the L2 regularizer utilised in our CBAM VGG16 is given in Eq. 10.
@@ 3.3 Explainability Techniques
3.3.1 Attention Based Technique
The inherent CBAM [woo2018cbamconvolutionalblockattention] layers of the models leverages with attention mechanisms to focus on the most informative spatial and channel-wise features. Unlike gradient-based approaches, these methods generate attention maps inherently from the model’s structure, offering insight into where the model looks without relying on backpropagation. CBAM attention maps embeds attention mechanisms directly within the network architecture unlike the post-hoc interpretability methods like GradCAM and GradCAM++. CBAM sequentially applies two types of attention, channel attention and spatial attention to refine the intermediate feature maps during the forward pass. The attention maps from CBAM can also be visualized to interpret the model’s focus at various stages. After each MaxPooling layer, the CBAM module generates:
  • Channel attention map ($C_{m}\in\mathbb{R}^{C\times 1\times 1}$): These maps highlights the most informative feature channels.
  • Spatial attention map ($S_{m}\in\mathbb{R}^{1\times H\times W}$): These maps highlight important spatial locations within the feature maps.
These attention maps offer real-time interpretability during training and inference by highlighting where the network focuses at each layer.
3.3.2 Post-hoc Attribution Based Explainability Methods
In this section, we explore different post-hoc explainability techniques used to interpret the model’s decision-making process.
Class-Discriminative Attribution Technique
Class-discriminative methods, such as Grad-CAM [Selvaraju2019] and Grad-CAM++ [Chattopadhay2018], utilize the gradients of class scores with respect to intermediate feature maps to produce localization maps. Grad-CAM is a widely-used technique to show the areas of an input image that are distinctive to a class and have the biggest influence on a model’s judgment. Grad-CAM++ is a generalized version of Grad-CAM that enables more accurate localization, especially in scenarios with several instances of the same item. These techniques highlight areas of the input that have the most influence for predicting a specific class, offering visually interpretable, class-specific explanations.
Layer-wise Relevance Propagation (LRP)
LRP [Bach2015], [Montavon2017] belongs to the class of additive explanation techniques, which also includes Shapley values [Lundberg2017], Gradient $\times$ Input [Simonyan2014], and DeepLIFT [Shrikumar2019]. These methods operate under the assumption that a function $f_{j}$ with $N$ input features $x=\{x_{i}\}_{i=1}^{N}$ can be expressed as a sum of contributions from each input variable, represented as $R_{i\leftarrow j}$, referred to as relevance scores. Here, $R_{i\leftarrow j}$ quantifies how much the $i$th input contributes to the $j$-th output. The total function value can be approximated (or exactly recovered) as shown in Eq. 11.
@@ 3.3.2 Post-hoc Attribution Based Explainability Methods
These relevance scores are propagated from the output layer back through the network to the input layer by the LRP algorithm. This backward redistribution is proportional to the contribution of each neuron in the forward computation, effectively assigning an importance score to each input feature. Unlike other interpretability approaches, LRP explicitly treats the neural network as a hierarchical, layer-wise acyclic graph where each unit $j$ in layer $l$ is associated with a local function $f^{l}_{j}$. Relevance values from the output layer $L$, where $R^{L}_{j}\propto f^{L}_{j}$, are passed backward, layer by layer, through the network until it reaches the input. This reverse mapping follows the same activation path used during the forward inference process, moving from the output node $f^{L}$ down to the initial input $f^{1}$, as illustrated in Fig. 6.
Figure 6: Visualization of the LRP mechanism. Relevance score $R_{j}$ or $R_{k}$, denotes how much an individual neuron contributes to the final output prediction
A key characteristic of LRP is its relevance conservation property, shown in Eq. 13, which ensures that the total relevance remains consistent across all layers of the model. This makes the interpretation more robust, as it maintains a one-to-one correspondence with the original model output $f^{L}$.
$R^{l-1}=\sum_{i}R^{l-1}_{i}=\sum_{i,j}R^{(l-1,l)}_{i\leftarrow j}=\sum_{j}R^{l}_{j}=R^{l}$ (13)
guaranteeing that the total relevance across each layer stays the same enables interpretable attribution, as each relevance score can be directly associated with the model’s original output $f^{L}$.
To address limitations of standard LRP, several rule-based variants have been proposed that tailor the relevance redistribution rules to specific layer types and activation behaviors. These methods include the below variants of LRP.
i. Epsilon Plus Rule ($\varepsilon^{+}$)[Bach2015]: This variant stabilizes relevance propagation by adding a small term $\varepsilon$ in the denominator, preventing division by small values and suppressing noise in low-activation neurons as given in Eq. 14.
$R_{i}=\sum_{j}\frac{z_{ij}}{z_{j}+\epsilon\cdot\text{sign}(z_{j})}R_{j}$ (14)
where $z_{ij}=x_{i}w_{ij}$ is the contribution of input neuron $i$ to output neuron $j$, $z_{j}=\sum_{i}z_{ij}$ is the total input to neuron $j$, and $\text{sign}(z_{j})$ denotes the sign of $z_{j}$.
ii. Epsilon Plus Gamma Box Rule ($\varepsilon^{+}\gamma\boxdot$) [anders2023softwaredatasetwidexailocal]: This composite rule introduces both the $\varepsilon$-stabilizer and a $\gamma$-scaling factor that increases the relevance assigned to positively activated neurons, thereby enhancing the interpretability of positively contributing features as given in Eq. 15.
$R_{i}=\sum_{j}\left(\frac{z^{+}_{ij}}{z^{+}_{j}+\epsilon}+\gamma\cdot\frac{z^{-}_{ij}}{z^{-}_{j}+\epsilon}\right)R_{j}$ (15)
where $z^{+}_{ij}=\max(0,z_{ij})$ and $z^{-}_{ij}=\min(0,z_{ij})$ represent the positive and negative contributions respectively, $z^{+}_{j}=\sum_{i}z^{+}_{ij}$ and $z^{-}_{j}=\sum_{i}z^{-}_{ij}$, and $\gamma$ is a scaling parameter that amplifies the relevance of positive contributions.
iii. Epsilon Plus Flat Rule ($\varepsilon^{+}_{\text{flat}}$)[Montavon2017]: This flat variant evenly distributes relevance across all connected inputs after applying $\varepsilon$-stabilization. It is particularly useful in fully connected or linear layers where uniform contribution is desired as given in Eq. 16.
$R_{i}=\sum_{j}\frac{1}{N_{j}}R_{j}$ (16)
where $N_{j}$ is the number of input neurons connected to neuron $j$.
iv. Epsilon Alpha2 Beta 1 Flat Rule ($\varepsilon^{\alpha=2,\beta=1}_{\text{flat}}$)[anders2023softwaredatasetwidexailocal]: A more aggressive redistribution scheme combining $\varepsilon$-stabilization with a relevance split governed by $\alpha=2$ and $\beta=1$, prioritizing positive contributions and suppressing negative evidence. It is effective for visualizing fine-grained relevance in deeper layers and attention-enhanced models like CBAM-VGG16 as given in Eq. 17.
$R_{i}=\sum_{j}\left(\alpha\cdot\frac{z^{+}_{ij}}{z^{+}_{j}}-\beta\cdot\frac{z^{-}_{ij}}{z^{-}_{j}}\right)R_{j}$ (17)
where $\alpha$ and $\beta$ are relevance weighting parameters (with $\alpha=2$, $\beta=1$), and $z^{+}_{ij}$, $z^{-}_{ij}$, $z^{+}_{j}$, and $z^{-}_{j}$ are defined as in Eq. 23.
These rules enable a more flexible and stable relevance propagation pipeline, making them well-suited for diverse architectural components in convolutional neural networks.
@@ 4 Experimental Results
The outcomes of our proposed methodology are shown in this section, together with information on the dataset used, experimental setup, evaluation metrics, results, and explainability analysis.
Table 1: Description of datasets used for plant disease classification
@@
Name of Disease/Crops
PlantVillage 54,303 43,442 10,861 38 Apple Scab, Grape Esca, Tomato Early Blight, etc.
Embrapa 46,376 37,100 9,276 93 Coffee, Cotton, Soybean, Citrus, Grapevine etc.
Maize 3,852 3,082 770 4 Gray Leaf Spot, Northern Leaf Blight, and Common Rust.
Apple 3,644 2921 723 4 Cedar Apple Rust, Healthy, Apple Scab etc.
Rice 9,000 7,200 1,800 10 Brown Spot, Blast, Tungro etc.
4.1 Datasets
We have selected five different dataset for to evaluate our proposed method. These five datasets, namely, PlantVillage, Embrapa, Maize, Apple, and Rice ensures the generalizability of our approach in plant leaf disease detection for other crops. All the datasets are divided in almost 80:20 ration in the training and testing set. The overall composition of the datasets is provided in Table 1.
@@ 4.2 Evaluation Metrics
The F1 score (F1), Area Under Curve (AUC), Recall (REC), Accuracy (ACC), Precision (PREC), and Cohen’s skappa score (KAPPA) are common classification metrics that are used to assess and compare the performance of the proposed CBAM-VGG16 model with state-of-the-art (SOTA) models. The evaluation is conducted across five publicly available datasets to ensure robustness and generalizability. The model’s feature representation capability is further assessed through high-dimensional visualization techniques, t-SNE [JMLR:v9:vandermaaten08a] and UMAP [mcinnes2020umapuniformmanifoldapproximation].
Dataset Loss Acc Prec Rec F1 AUC Kappa
Apple 0.41 95.42 95.27 95.42 95.27 99.05 0.90
Embrapa 0.32 94.20 94.70 93.90 94.30 99.75 0.95
Maize 0.12 95.00 95.05 95.00 95.01 99.62 0.93
PlantVillage 0.03 98.72 98.72 98.72 98.72 99.95 0.99
Rice 0.01 98.87 98.87 98.87 98.87 99.94 0.99
Table 2: Performance of CBAM-VGG16 model on different datasets
@@
Table 3: Comparison of the work with other SOTA methods on five publicly available datasets
4.3 Results and Discussion
The evaluation findings of the suggested CBAM-VGG16 model on five publicly accessible plant disease datasets—Apple, Embrapa, Maize, PlantVillage, and Rice—are shown within this section. According to Table 2, the model demonstrates strong generalization capabilities, achieving high accuracy across all datasets, with the best performance observed on the Rice (98.87%) and PlantVillage (98.72%) datasets. These accuracy scores are complemented by consistently high values of precision, recall, F1-score, and AUC, indicating balanced and reliable classification performance across classes. In addition, the model’s effectiveness is further supported by Cohen’s Kappa statistics, which range from 0.90 to 0.99. Notably, the highest Kappa values (0.99) on the Rice and PlantVillage datasets show almost perfect consistency between the actual and predicted labels, outperforming chance and confirming the validity of the suggested approach. Together, these findings show how reliable and successful the CBAM-VGG16 model is at identifying plant diseases in a variety of datasets.
Additionally, we assess our model’s performance by comparing it with recent techniques. The comparative performance is summarized in Table 3. It can be observed from the results that across all five datasets, our CBAM-VGG16 consistently outperforms existing models [Karthik2020], [CHEN2021], [Chen2021rice], [THAKUR2023], [Hangli], [chenjunde], [YANG2023107543], [PSTHAKUR], [li2024yololeaf], [shafik2024transfer], [Bhuyan2024], [Alpsalaz2025]. On the Apple dataset, the our method performs better with the highest accuracy of 95.42%. In other evaluation metrics also our proposed method demonstrates superior precision, recall, and F1-score, along with an AUC of 99.05% and a Cohen’s kappa score of 0.90. On the Embrapa dataset, CBAM-VGG16 again achieves best performance, with an accuracy of 94.20%. The similar superiority performance is observed in other evaluation metrics as well, indicating the robust generalization capability of our proposed method across different datasets. For the Maize dataset, our model obtains a notable accuracy of 95.00%, outperforming all the methods while also achieving the best F1-score, AUC (99.62%), and Kappa score (0.93). In the PlantVillage dataset, the proposed model maintains a competitive edge with nearly perfect scores across all metrics. Notably, it again exceeds in performance, reaffirming its superior feature learning and attention capabilities as compared with other methods. Lastly, on the Rice dataset, our approach attains the maximum accuracy of 98.87%, alongside the best precision, recall, and AUC (99.94%), demonstrating its effectiveness in capturing fine-grained disease patterns even in challenging samples. The obtained results on all the datasets exhibits that the suggested approach has good generalizability and can be adapted to identify leaf diseases in any other crop.
Figure 7: Illustrations of CBAM attention maps generated for inputs of different classes on a CBAM-VGG16. Rows from top to bottom corresponds to images of class black rot, esca, leaf blight and healthy, respectively.
Figure 8: Visualization of GradCAM, GradCAM++ outputs. Rows from top to bottom corresponds to images of class black rot, esca, leaf blight and healthy, respectively.
@@ 4.4 Explainability Analysis
We have selected grape leaf dataset from plant village dataset for detailed explainability analysis of our achieved results to validate the model’s interpretability.
@@
Figure 10: Explainability analysis for other crop leafs used in the experiments. Top to bottom row represents leafs of Apple, Embrapa (Coconut), Maize, and Rice.
4.4.1 CBAM Attention Maps Visualization
In the Fig. 7, CBAM attention maps generated across all the five convolutional stages inserted in the proposed CBAM-VGG16 model for grape leaf classification are shown. These maps provide insight into how the model attends to informative regions at various depths of the network. It can be seen that the first CBAM block (CBAM1) provides an attention maps highlighting low-level features such as leaf textures, edges, and background suppression. These early filters focus on enhancing basic visual cues while discarding irrelevant background. The subsequent CBAM blocks, further refines these maps where focus is diverted to distinguishable disease prone areas in the leaf. Overall, these visualizations demonstrate how CBAM refines feature learning progressively across the network, resulting in improved focus, interpretability, and diagnostic accuracy.
4.4.2 Attribution Based Explainability Methods
To better understand the internal reasoning of the CBAM-VGG16 model predictions, we examine and compare visual explanations generated using, Grad-CAM, and Grad-CAM++ (post-hoc explainability techniques). Fig. 8 displays a visual comparison of these heatmaps for various leaf samples across different disease classes. Grad-CAM computes gradients reaching the last convolutional layers to determine which input locations most affected the model’s decision-making. However, it typically produces relatively coarse, blob-like heatmaps. In contrast, Grad-CAM++ refines this by incorporating second-order gradients, resulting in more spatially accurate and class-discriminative attention maps. Despite their usefulness, both are post-hoc techniques, which do not influence or enhance the training process directly.
4.4.3 Layerwise Relevance Propagation (LRP)
Fig. 9 illustrates the class-wise attribution heatmaps generated for the CBAM-VGG16 model using a comprehensive suite of LRP techniques. The columns represent different attribution methods of LRP ,namely Epsilon Gamma Box, Epsilon Plus, Epsilon Plus Flat, and Epsilon Alpha2 Beta1 Flat. The LRP family consistently produces sparse yet discriminative saliency maps, aligning well with regions exhibiting disease-related symptoms. Among them, Epsilon Plus Flat and Epsilon Alpha2 Beta1 Flat stand out for their capacity to sharply localize high-relevance areas, typically corresponding to lesions, necrotic margins, or discolored patches symptomatic of disease. These variants effectively suppress irrelevant background and vein structures, enhancing focus on pathologically significant textures.
We have shown in Fig. 10 the explainable visualization of proposed method on some other disease of other dataset used in this work. These visualization also exhibits the same properties as observed in the detailed explainability analysis for grape leaf dataset.

### v2

@@ 3.1.2 Spatial Attention Module
This module generates a 2D spatial attention map $S_{m}\in\mathbb{R}^{c\times W\times H}$, which is applied to the channel-refined feature map $I_{c}f^{\prime}$. To compute spatial attention, the module first applies average pooling and max pooling operations along the channel axis, resulting in two 2D feature maps: $I_{s}^{\text{avg}}\in\mathbb{R}^{1\times W\times H}$ and $I_{s}^{\text{max}}\in\mathbb{R}^{1\times W\times H}$. These two maps are then concatenated along the channel dimension and passed through a convolutional layer to produce the spatial attention map. The resulting map is used to refine the features by performing an element-wise multiplication with $I_{c}f^{\prime}$, yielding the final spatially-refined output $I_{s}f^{\prime\prime}$. The spatial attention computation in CBAM is formally described in Eqs. 5 and  6.
$\footnotesize S_{m}(I_{c}f^{\prime})$ $=\sigma\big(\nu_{7\times 7}([\text{AvgPool}(I_{c}f^{\prime});\text{MaxPool}(I_{c}f^{\prime})])\big)$ (5)
$\footnotesize S_{m}(I_{c}f^{\prime})$ $=\sigma\big(\nu_{7\times 7}([I_{s}^{\text{avg}};I_{s}^{\text{max}}])\big)$ (6)
In the above equations, $\sigma$ represents the sigmoid activation function, and $\nu_{7\times 7}$ denotes a convolution operation with a kernel size of $7\times 7$.
The intermediate layers of the model utilize the Rectified Linear Unit (ReLU) activation function, defined as given in Eq. 7.
@@
The final output layer employs the Softmax activation function, a normalized exponential function commonly used for multi-class classification. Softmax transforms the input vector into a probability distribution where the sum of all output probabilities is one, as defined in Eq. 8.
@@
In our work, as indicated by Eq. 9, we have utilised the Cross Entropy loss function, for model optimization.
@@
To reduce the model’s complexity, L2 regularization is employed where extreme changes in weights during the training phase are penalized by regularization reducing the probability of over-fitting. During training, the L2 regularizer additionally simplifies the input features and stabilises performance. The definition of the L2 regularizer utilised in our CBAM VGG16 is given in Eq. 10.
@@ 3.2 Explainability Techniques
3.2.1 Post-hoc Attribution Methods
Methods such as Grad-CAM [Selvaraju2019] and Grad-CAM++ [Chattopadhay2018], utilize the gradients of class scores with respect to intermediate feature maps to produce localization maps. Grad-CAM is a widely-used technique to show the areas of an input image that are distinctive to a class and have the biggest influence on a model’s judgment. Grad-CAM++ is a generalized version of Grad-CAM that enables more accurate localization, especially in scenarios with several instances of the same item.
3.2.2 LRP Techniques
LRP [Bach2015], [Montavon2017] belongs to the class of additive explanation techniques. These methods operate under the assumption that a function $f_{j}$ with $N$ input features $x=\{x_{i}\}_{i=1}^{N}$ can be expressed as a sum of contributions from each input variable, represented as $R_{i\leftarrow j}$, referred to as relevance scores. Here, $R_{i\leftarrow j}$ quantifies how much the $i$th input contributes to the $j$-th output. The total function value can be approximated (or exactly recovered) as shown in Eq. 11.
@@ 3.2.2 LRP Techniques
Unlike other interpretability approaches, LRP explicitly treats the neural network as a hierarchical, layer-wise acyclic graph where each unit $j$ in layer $l$ is associated with a local function $f^{l}_{j}$. Relevance values from the output layer $L$, where $R^{L}_{j}\propto f^{L}_{j}$, are passed backward, layer by layer, through the network until it reaches the input. This reverse mapping follows the same activation path used during the forward inference process, moving from the output node $f^{L}$ down to the initial input $f^{1}$.
We have also used some of the advanced variants of LRP such as Epsilon Plus Rule ($\varepsilon^{+}$)[Bach2015], Epsilon Plus Gamma Box Rule ($\varepsilon^{+}\gamma\boxdot$) [anders2023softwaredatasetwidexailocal], Epsilon Plus Flat Rule ($\varepsilon^{+}_{\text{flat}}$)[Montavon2017], and Epsilon Alpha2 Beta 1 Flat Rule ($\varepsilon^{\alpha=2,\beta=1}_{\text{flat}}$)[anders2023softwaredatasetwidexailocal] for assesing the interpretability of our proposed method.
@@ 4 Experimental Results
4.1 Datasets
We have selected five different datasets for to evaluate our proposed method. These five datasets, namely, PlantVillage, Embrapa, Maize, Apple, and Rice ensure the generalizability of our approach in plant leaf disease detection for other crops. All the datasets are divided in an almost 80:20 ratio in the training and testing set. The overall composition of the datasets is provided in Table 1.
Table 1: Description of different datasets used in our work.
@@ 4.1 Datasets
Disease/Crops
PlantVillage 54,303 43,442 10,861 38
Apple Scab, Grape
Esca, Tomato Early
Blight, etc.
Embrapa 46,376 37,100 9,276 93
Coffee, Cotton,
Soybean, Citrus,
Grapevine etc.
Maize 3,852 3,082 770 4
Gray Leaf Spot,
Northern Leaf
Blight etc.
Apple 3,644 2921 723 4
Cedar Apple Rust,
Healthy, Apple Scab etc.
Rice 9,000 7,200 1,800 10
Brown Spot, Blast,
Tungro etc.
@@ 4.2 Evaluation Metrics
The F1 score (F1), Area Under Curve (AUC), Recall (REC), Accuracy (ACC), Precision (PREC), and Cohen’s Kappa score (KAPPA) are common classification metrics that are used to assess and compare the performance of the proposed CBAM-VGG16 model with state-of-the-art models. The model’s feature representation capability is further assessed through high-dimensional visualization techniques, t-SNE [JMLR:v9:vandermaaten08a] and UMAP [mcinnes2020umapuniformmanifoldapproximation].
Table 2: Comparison of the work with other state-of-the-art methods on five publicly available datasets.
@@
4.3 Results and Comparative Study
The performance results along with a detailed comparative study is reported in Table 2. It can be observed from the results that across all five datasets, our CBAM-VGG16 consistently outperforms existing models [Karthik2020], [CHEN2021], [Chen2021rice], [THAKUR2023], [Hangli], [chenjunde], [YANG2023107543], [PSTHAKUR], [li2024yololeaf], [shafik2024transfer], [Bhuyan2024], [Alpsalaz2025]. On the Apple dataset, the our method performs better with the highest accuracy of 95.42%. On the Embrapa dataset, CBAM-VGG16 again achieves best performance, with an accuracy of 94.20%. The similar superiority performance is observed in other evaluation metrics as well, indicating the robust generalization capability of our proposed method across different datasets. For the Maize dataset, our model obtains a notable accuracy of 95.00%, outperforming all the methods. In the PlantVillage dataset, the proposed model maintains a competitive edge with nearly perfect scores across all metrics. Notably, it again exceeds in performance, reaffirming its superior feature learning and attention capabilities as compared with other methods. Lastly, on the Rice dataset, our approach attains the maximum accuracy of 98.87%, alongside the best precision, recall, and AUC (99.94%), demonstrating its effectiveness in capturing fine-grained disease patterns even in challenging samples. The obtained results on all the datasets exhibits that the suggested approach has good generalizability and can be adapted to identify leaf diseases in any other crop.
Figure 2: Illustrations of CBAM attention maps generated for inputs of different classes on a CBAM-VGG16. Rows from top to bottom correspond to images of class black rot, esca, leaf blight, and healthy, respectively.
Figure 3: Visualization of GradCAM and GradCAM++ outputs. Rows from top to bottom correspond to images of class black rot, esca, leaf blight, and healthy, respectively.
@@ 4.4 Explainability Analysis
Figure 5: Explainability analysis for other crop leaves used in the experiments. Top to bottom row represents leaves of Apple, Embrapa (Coconut), Maize, and Rice.
We have selected grape leaf dataset from plant village dataset for detailed explainability analysis of our achieved results to validate the model’s interpretability. In the Fig. 2, CBAM attention maps generated across all the five convolutional stages inserted in the proposed CBAM-VGG16 model for grape leaf classification are shown. These maps provide insight into how the model attends to informative regions at various depths of the network. Overall, these visualisations demonstrate how CBAM refines feature learning progressively across the network, resulting in improved focus, interpretability, and diagnostic accuracy.
Fig. 3 displays a visual comparison of Grad-CAM and Grad-CAM++ heat maps for various leaf samples across different disease classes. Grad-CAM computes gradients reaching the last convolutional layers to determine which input locations most affected the model’s decision-making. However, it typically produces relatively coarse, blob-like heatmaps. In contrast, Grad-CAM++ refines this by incorporating second-order gradients, resulting in more spatially accurate and class-discriminative attention maps.
Fig. 4 illustrates the class-wise attribution heatmaps generated for the CBAM-VGG16 model using a comprehensive suite of LRP techniques. The LRP family consistently produces sparse yet discriminative saliency maps, aligning well with regions exhibiting disease-related symptoms. Among them, Epsilon Plus Flat and Epsilon Alpha2 Beta1 Flat stand out for their capacity to sharply localize high-relevance areas, typically corresponding to lesions, necrotic margins, or discolored patches symptomatic of disease. These variants effectively suppress irrelevant background and vein structures, enhancing focus on pathologically significant textures.
We have shown in Fig. 5 the explainable visualization of proposed method on some other disease of other dataset used in this work. These visualisations also exhibit the same properties as observed in the detailed explainability analysis for grape leaf dataset.